John has 6 cookies and 3 protein bars, while Luke has 3 cookies and 6 protein bars. Although they like the same products, they have different preferences over them; John’s utility function is ?(?,?)= ??2 and Luke’s utility function is ?(?,?)=?2? . The two decide to engage in trade.

a. Without normalizing prices, find their demand functions for each good.

b. Write the excess demand function for cookies.

c. Normalizing the price of protein bars, find the equilibrium price of cookies (??∗) that clears the two markets. How many cookies and protein bars does each of them consume in the competitive equilibrium?

d. Are they each better off after engaging in trade than they would have been consuming their endowments? Show your calculations.
Answer #1

web ; wid=3 J represent juhs (superscripts) .. o cookiess; b bars Coubscripts) w = 3; wil 6= 6 L represento luke (supersult.)substitute in budget constraint & Pet *ъ (2.с) - fro+ 3 г. Pe c+2ce = c р.+ 3r. 3c [ .) - Cro+ 3rъ Johns [= opet 3p6 demand- Excen demand for cookies, der 1 demand for cookies demand of wokies - supply of corkies | Lab - ct - (- ] =ope + 3 p +2 p.no (19= 3; 65-6 Joon congumes 3 lukies P & 6 barn 3. Luke 6 cookies consumes 1 3 boen ability of tor bade Johns utility: ve